A shop[ sells 5l water bottles.
a)On Wednesday the shop received $3450 from selling5l bottles of water at $11.50.How many bottles were sold?
b)On Thursday,the shop received $x by selling bottles of water at $10each.In
terms of x,how many bottles of water were sold?
c)On Friday the shop received $(x+10) by selling bottles of water at $12each.In terms of x,how many bottles of water were sold?

Answer by karaoz(32)


I wonder who is buying all those water bottles. What country is this? Rather bizarre thing ... this... buying water! Selling, on the other hand, is rather understandable since people will try to sell anything anyway. This particular shop seems to be doing rather well.

Let's see about Wednesday.
They received $3450 (Wow!) from selling 5l bottles of water at $11.50 (Wow! Wow! Auch!). How many bottles were sold?
Since we need to know how many bottles were sold, this is a reasonable definition for the unknown value (and we will name it by using letter b). So,
b = "the number of bottles sold".
In short, the question is then
$11.50 * b = $3450. What is b?
b must be $3450 / $11.50 = 300.

How about Thursday?
The revenue was $x and the price per bottle was $10. What we need to find out is again the same thing, the number of bottles sold, b.
$10 * b = $x. What is b?
b must be $x/$10, or simply b = x/10.

Friday must be the best of all. People go crazy and buy 5l bottles of water for $12!
OK, similar story, similar question.
This time the price per bottle is $12 and the revenue is $(x+10). So,
$12 * b = $(x+10). What is b?
b = $(x+10)/$12, or simply (x+10)/12.

Hope this helps!